# Context Matters: Peer-Aware Student Behavioral Engagement Measurement via VLM Action Parsing and LLM Sequence Classification

### v1

@@ Abstract
Understanding student behavior in the classroom is essential to improve both pedagogical quality and student engagement. Existing methods for predicting student engagement typically require substantial annotated data to model the diversity of student behaviors, yet privacy concerns often restrict researchers to their own proprietary datasets. Moreover, the classroom context, represented in peers’ actions, is ignored. To address the aforementioned limitation, we propose a novel three-stage framework for video-based student engagement measurement. First, we explore the few-shot adaptation of the vision-language model for student action recognition, which is fine-tuned to distinguish among action categories with a few training samples. Second, to handle continuous and unpredictable student actions, we utilize the sliding temporal window technique to divide each student’s 2-minute-long video into non-overlapping segments. Each segment is assigned an action category via the fine-tuned VLM model, generating a sequence of action predictions. Finally, we leverage the large language model to classify this entire sequence of actions, together with the classroom context, as belonging to an engaged or disengaged student. The experimental results demonstrate the effectiveness of the proposed approach in identifying student engagement.
@@ 1 Introduction
Students’ active participation and engagement in the classroom are crucial for successful learning outcomes [13]. Student engagement is a multidimensional construct that encompasses behavioral, emotional and cognitive engagement [17]. Behavioral engagement refers to students’ self-directed actions to get access to the curriculum, such as studying and doing homework, hand and body movements while attending lectures, as well as engaging cooperatively in classroom activities [4]. Emotional engagement, on the other hand, refers to students’ feelings about their learning experience, teachers, and peers [27]. This involves feelings of happiness in learning, disinterest in the subject matter, or frustration and struggle to understand. Lastly, cognitive engagement is the psychological investment in academic achievement, which is usually revealed by a student’s ability to comprehend, organize, and analyze in the act of deep learning to fully understand the material [10].
@@
Behavioral engagement is a key indicator of student participation and attentiveness, with a direct impact on academic performance [5]. During a lecture, a student’s behavior can reveal whether they are on- or off-task. While a disengaged student engages in off-task activities like playing with a mobile device and checking the time, an engaged student typically listens, takes notes, and types on a laptop. Traditional assessments of this engagement depend on manual classroom observations [28, 19]. While informative, these observations are labor-intensive and subject to observer bias, which limits their scalability in large educational settings. A key aspect of engagement annotation is its temporal resolution, which defines the time scale at which engagement is evaluated. Following the standard classroom observation protocols [28, 19], we adopt a 2-minute time resolution, i.e., the length of each video sample is set to match a 2-minute time window.
@@
Automating student behavioral engagement in classrooms has received considerable attention due to advances in computer vision and deep learning. Existing studies for student behavior analysis can be categorized according to the target outcome into action recognition, learning analytics, and engagement measurement. Methods for recognizing student actions focus on identifying a particular group of students’ behaviors without conducting additional analysis, providing instructors with minimal useful information. Furthermore, a considerable number of studies [20, 36] formulated student behavior analysis as single- or multi-object detection, ignoring the complexity and diversity of student behaviors. Learning analytics approaches, which use current computer vision techniques to analyze student behavior, provide a set of visual features but little useful information for instructors. In these approaches, a specific student’s actions, gaze, and/or facial expression are usually employed in estimating the student’s engagement.
@@
Unlike traditional Human Action Recognition (HAR) models, which were trained on clips that have been trimmed to include only one action, detecting student actions over arbitrary time intervals necessitates the analysis of continuous, untrimmed video streams in which the execution and order of actions are critical. To bridge this gap, research has shifted toward methods for temporal action localization (TAL) [29, 34], which identify and locate the start and end times of sparse, intermittent actions within long, untrimmed videos. However, TAL tends to ignore “background” segments (e.g., listening, looking to the side/back), which are important in engagement analysis. Moreover, Yu et al., [34] leveraged LLM to generate a behavior analysis report without estimating students’ engagement. In order to measure students’ behavioral engagement, Abdelkawy et al. [2] created a histogram of the students’ actions over an arbitrary time period. The video clip was divided into overlapping segments using the sliding temporal window technique over a 2-minute period. These segments are then fed into a 3D CNN model for action recognition. A classifier that assesses the student’s level of engagement uses both this histogram and the student’s gaze as inputs. However, the student’s actions throughout such a time interval are not manually annotated. Thus, the evaluation of the behavioral engagement classifier remains incomplete.
@@
The limitations of the three types of approaches are:
  • There is a scarcity in publicly available datasets for students actions and engagement, since in each studey, the authors collect their proprietary dataset due to privacy concerns.
  • The context, which is defined as actions performed by peers while the student of interest is performing a certain action, is underestimated or ignored (see Fig. 1).
Since student engagement depends not only on their actions but also on the current classroom activity and the behavior of their peers, classroom context is crucial to accurately assessing student engagement. This is especially important as contemporary pedagogy increasingly moves beyond passive lecture-based learning. ICAP framework [10] characterizes learning activities along a Passive–Active–Constructive–Interactive hierarchy, where interactive tasks generally reflect deeper cognitive engagement. The same observable behavior can therefore signal opposite engagement states depending on context. A student turning to talk to a peer during a traditional lecture, for instance, probably shows disengagement, but doing the same during a peer-discussion activity that is prompted shows engagement. When interpreting any student action in a classroom, context is important. For example, when a student asks a question, other students are expected to turn to listen to their peer; when a student is asked to work independently on a problem or possibly with a partner, they are engaged by focusing intently on their laptop for problem solving; when the instructor is lecturing or demonstrating a problem solution, on the other hand, they are disengaged by focusing on their laptop while accessing their social media feeds.
We propose a novel three-stage framework for classifying behavioral engagement based on student body movement in order to overcome the aforementioned drawbacks. In order to classify human actions in trimmed videos with a limited number of labeled samples per category, we first perform few-shot adaptation for the vision-language model [23]. Second, because the student actions are continuous, we divide each 2-minute time-resolution student video into non-overlapping segments using the sliding temporal window technique. The VLM model (from stage 1) is then used to assign a predefined label to each segment, resulting in a time-ordered sequence of predicted actions. Third, an LLM uses this sequence of actions and the classroom context as input to determine whether a student is engaged or disengaged.
To the best of our knowledge, prior studies have not investigated whether current vision-language models (VLMs) and large language models (LLMs) can effectively recognize the student’s class actions and measure behavioral engagement based on identified actions, respectively. Accordingly, this paper addresses the research questions:
@@
    Compared to traditional action classification, is it harder for VLMs to classify actions of students in-class?
@@
    Which action sequence or action histogram most accurately depicts students’ behavior?
@@
  • Develop a novel framework for measuring student behavioral engagement, which integrates VLM for action recognition and LLM for engagement classification. To address the lack of publicly available datasets for student actions and engagement, we perform a few-shot tuning for VLM to improve student action recognition, and we use the LLM in a zero-shot prompting for engagement classification.
  • Create a novel student behavior dataset with two components: 1) trimmed videos (13 categories, controlled environment) with annotated students’ actions; and 2) untrimmed videos (3 lectures, 11 students) with annotated students’ engagement. The experts’ annotations of this dataset include ”engaged/disengaged” labels for each 2-minute video, and dense temporal action labels for each segment within them.
@@ 2 Related Work
2.1 Measuring Student Engagement
@@
The emerging and promising real-time engagement measurement approach, enabled by advances in computer vision and machine learning, allows tracking a student’s engagement at fine-grained temporal resolutions on the order of seconds or minutes through discrete, objective indicators. Existing research of student behavior analysis in classrooms can be grouped into three categories: student action recognition, learning analytics, and engagement measurement. While student action recognition approaches focus on the student’s actions from trimmed video or a single image, learning analytics approaches rely on facial expressions, student actions, and/or head gaze. For example, Sheng et al. [25] proposed an enhanced YOLOv8-based detector for localizing and classifying student behaviors from single classroom images. Their framework incorporates a Multi-Scale Large Kernel Convolution Module (MLKCM) to extract multiscale spatial features and a Progressive Feature Optimization Module (PFOM) that segments the channel dimension of the input feature map for improved feature refinement. Ahuja et al. [3] presented a classroom sensing system (EduSense) that provides a set of features, including detection of hand raise, hand on face, cross arms, smile, head orientation, and speech duration for both the teacher and the students. The authors employed students’ skeleton joints and students’ facial landmarks to extract these visual features due to the presence of much visual noise, occlusion in the classroom, and privacy preservation. The student’s actions are recognized frame-wise, which leads to a high false-positive rate in recognition, especially for hand raising. Line et al.[21] recognized student behavior in the classroom based on upper body skeleton joints. Their framework includes three steps: Extracting students’ poses using bottom-up pose estimation (Openpose) and removing the incorrect connection of students’ joints using YoloV4 object detection, generating feature vector using normalized joint locations, joint distances, and bone angles to represent student posture, and using four fully connected layer neural network to classify the student feature vector into four classes. The main limitations of their framework are that it recognizes student actions from single images rather than video segments, and it considers actions with different engagement levels, such as writing and playing with a phone as one class (bowing).
@@
Student engagement studies, which employ student actions and/or head gaze to measure the student behavioral engagement, provide the teachers with actionable insights. In [35], the authors proposed a framework that processes classroom videos to first estimate each student’s 2D pose per frame and then apply a Temporal Action Detection (TAD) [16] module to produce a temporal sequence of actions for each student. These per-student action sequences are subsequently fed into an LLM to generate behavioral reports. The TAD module was trained on the NTU RGB+D dataset [24], using only action classes relevant to classroom contexts. However, this approach has two main limitations. First, it assumes that each student’s behavior is independent of others, disregarding the broader classroom context. Second, the model does not predict explicit engagement labels; instead, it outputs textual behavioral reports that still require human interpretation to extract meaningful insights. Abdelkawy el al.[2] utilized a random forest classifier to detect whether the student is on-task or off-task based on a histogram of actions, together with head gaze. Each student’s 2-minute-long video was divided into overlapping segments using the sliding temporal window technique; the student skeleton sequence of each segment is assigned a predefined action label using 3D CNN model and then the histogram is created from these identified actions. Nevertheless, the student’s actions during the given period are not manually annotated. Consequently, the evaluation of the behavioral engagement classifier is still incomplete. Unlike these studies, we analyze the entire 2-minute video to get the student’s sequence of actions, which are used for engagement measurement. Moreover, we constructed a dataset that has dual expert annotations: engaged/disengaged for each 2-minute video and its corresponding frame-wise action labels to measure the performance of the engagement classifier. We also investigated the few-shot adaptation of the vision-language model for student action recognition, which learns to differentiate between action classes with a few training samples to overcome the challenge of the scarcity of annotated data required to model the diversity of student behaviors.
@@ 2.3 Large Language Models
Large language models (LLMs) are Transformer-based models with hundreds of billions of parameters, trained on large-scale text corpora (e.g., GPT-3 [8], LLaMA [12], Gemma [30], etc.). They exhibit generalization abilities, enabling them to perform tasks in zero-shot settings without requiring task-specific training examples. While several works employed LLMs in education for multiple-choice question generation and assignment grading, there is less work that leverages the reasoning capability of LLMs in student behavior analysis. Yu et al. [34] utilized an LLM to generate a pedagogical report based on the structured temporal action information for each student, which includes action categories and their timestamps. In this work, we use the LLM to assign an engagement level to each student’s sequence of actions.
@@ 3.1 Overview of the proposed method
The proposed framework consists of three main steps: segment-wise action recognition, VLM action parsing, and engagement classification (Fig. 2). First, a vision-language model is adapted in a few-shot manner using a small set of labeled trimmed student action clips to improve recognition on continuous student videos. Second, the student’s untrimmed video is split into non-overlapping segments using a sliding-window technique, and the fine-tuned VLM assigns an action label to each segment. Consecutive identical predictions are merged to form an ordered action sequence that preserves both action identity and duration. Finally, this action sequence, along with the classroom context, is provided to an LLM to determine whether the student is engaged or disengaged. Representing student’s behavior during a 2-minute interval as a temporal sequence rather than a histogram, where the duration of each action is accumulated, allows the model to capture not only how long each action occurs but also their ordering, which is shown to be crucial for interpreting engagement. For example, a student may still be deemed engaged if they take notes for the majority of the time period but play on their phone for one continuous 25-second block (e.g., at the beginning or end). However, a student is considered disengaged if they alternate between taking notes and making frequent quick phone checks because this pattern of frequent disruption shows an inability to maintain cognitive focus on the task. Therefore, the temporal sequence and context of actions more accurately reflect engagement than the total amount of time spent in each action (e.g., via a histogram).
Furthermore, when representing student action as a sequence, together with peers’ action sequences, the LLM predictions are easier to be interpreted. Fig.3 shows examples of students’ actions aligned with classroom actions for engaged (a) and disengaged (b, c) students.
@@ 3.2 Actions dictionary
Since this study is conducted on college students in introductory engineering STEM classes, the actions they perform during classroom time, from which student behavior engagement can be estimated, are limited. We adopted classroom observation protocols in [2], and we came up with the following 13 actions:
  • Eating meal/snack,
  • Writing on notebook/tablet,
  • Typing on a laptop,
  • Playing with mobile phone,
  • Looking to the side/back,
  • Looking down w/o reading/writing,
  • Looking at laptop screen (not typing).
  • Raising hand,
  • Rubbing face,
  • Reading,
  • Drinking,
  • Yawning,
  • Listening,
Fig. 4 shows samples of such actions.

### v2

@@ Abstract
Understanding student behavior in the classroom is essential to improve both pedagogical quality and student engagement. Existing methods for predicting student engagement typically require substantial annotated data to model the diversity of student behaviors, yet privacy concerns often restrict researchers to their own proprietary datasets. Moreover, the classroom context, represented in peers’ actions, is ignored. To address the aforementioned limitation, we propose a novel three-stage framework for video-based student engagement measurement. First, we explore the few-shot adaptation of the vision-language model for student action recognition, which is fine-tuned to distinguish among action categories with a few training samples. Second, to handle continuous and unpredictable student actions, we utilize the sliding temporal window technique to divide each student’s 2-minute-long video into non-overlapping segments. Each segment is assigned an action category via the fine-tuned VLM model, generating a sequence of action predictions. Finally, we leverage the large language model to classify this entire sequence of actions, together with the classroom context, as belonging to an engaged or disengaged student. The experimental results demonstrate the effectiveness of the proposed approach in identifying student engagement. The source code and dataset will be available upon request.
@@ 1 Introduction
Students’ active participation and engagement in the classroom are essential for successful learning outcomes [30, 14]. Student engagement is a multidimensional construct comprising behavioral, emotional, and cognitive engagement [18]. Behavioral engagement refers to students’ self-directed actions to get access to the curriculum, such as studying and doing homework, hand and body movements while attending lectures, as well as engaging cooperatively in classroom activities [4]. Emotional engagement, on the other hand, refers to students’ feelings about their learning experience, teachers, and peers [28]. This involves feelings of happiness in learning, disinterest in the subject matter, or frustration and struggle to understand. Lastly, cognitive engagement is the psychological investment in academic achievement, which is usually revealed by a student’s ability to comprehend, organize, and analyze in the act of deep learning to fully understand the material [11, 8].
@@
Behavioral engagement is a key indicator of student participation and attentiveness, with a direct impact on academic performance [5]. During a lecture, students’ behaviors often reveal whether they are on- or off-task. While a disengaged student engages in off-task activities like playing with a mobile device and frequently checking the time, an engaged student typically listens, takes notes, and types on a laptop. Conventional assessments of behavioral engagement typically depend on manual classroom observations [29, 20]. While informative, these observations are labor-intensive and subject to observer bias, which limits their scalability in large educational settings. An important consideration in engagement annotation is temporal resolution, which defines the time scale over which engagement is assessed. Following the standard classroom observation protocols [29, 20], we adopt a 2-minute temporal resolution.
@@
Automating student behavioral engagement in classrooms has received considerable attention due to advances in computer vision and deep learning. Existing studies for student behavior analysis can generally be categorized according to the target outcome into three groups: action recognition, learning analytics, and engagement measurement. Action recognition approaches focus on identifying a particular group of students’ behaviors without conducting additional analysis, providing instructors with minimal useful information. Furthermore, a considerable number of studies [21, 38] formulated student behavior analysis as single- or multi-object detection, ignoring the complexity and diversity of student behaviors. Learning analytics approaches [3], which use current computer vision techniques to get a specific student’s actions, gaze, and/or facial expression, provide a set of visual features but little useful information for instructors.
@@
Unlike traditional Human Action Recognition (HAR) models that operate on single-action trimmed clips, student behavior analysis demands processing continuous, untrimmed videos in which the execution and order of actions are critical. To address this challenge, research has shifted toward temporal action localization (TAL) [31, 37], which identifies and locates the start and end times of sparse, intermittent actions within long, untrimmed videos. However, TAL tends to ignore “background” segments (e.g., listening, looking to the side/back), which are important in engagement analysis. Furthermore, while Yu et al., [37] leveraged a large language model (LLM) to generate a behavior analysis report, their approach does not explicitly estimate student engagement levels. To measure students’ behavioral engagement, Abdelkawy et al. [2] represented a student’s actions and their frequencies over an arbitrary time period as a histogram of actions. This histogram, along with the student’s gaze, is then used as input to an engagement classifier. However, the student’s actions within this time interval are not manually annotated, leaving the evaluation of the behavioral engagement classifier incomplete.
@@
Overall, existing approaches face two key limitations:
  • Scarcity of publicly available datasets for students actions and engagement, as most studies rely on proprietary datasets due to privacy concerns.
  • limited consideration of classroom context, which is defined as actions performed by peers while the student of interest is performing a certain action (see Fig. 1).
Student engagement depends not only on their actions but also on the ongoing classroom activity and the behavior of their peers; therefore, classroom context is crucial for accurate engagement assessment. This is particularly important as contemporary pedagogy increasingly moves beyond passive lecture-based learning. The ICAP framework [11] characterizes learning activities along a Passive–Active–Constructive–Interactive hierarchy, where interactive tasks generally reflect deeper cognitive engagement. The same observable behavior can therefore signal opposite engagement states depending on context. For example, a student turning to talk to a peer during a traditional lecture probably shows disengagement, but doing the same during a peer discussion activity that is prompted shows engagement. Similarly, when a student is asked to work independently on a problem, he is engaged by focusing intently on his laptop for problem-solving. In contrast, he is disengaged if he uses his laptop for unrelated activities such as browsing social media while the instructor is lecturing.
To address the aforementioned limitations, we propose a novel three-stage framework for classifying behavioral engagement based on student actions. First, we perform few-shot adaptation for the vision-language model (VLM)  [24] to classify human actions in trimmed videos with a limited number of labeled samples per category. Second, because the student actions are continuous, we divide each 2-minute time-resolution student video into non-overlapping segments using the sliding temporal window technique. The VLM model (from stage 1) is then used to assign a predefined label to each segment, resulting in a time-ordered sequence of predicted actions. Third, an LLM uses this sequence of actions and the classroom context as input to determine whether a student is engaged or disengaged.
To the best of our knowledge, prior studies have not investigated whether current vision-language models (VLMs) and large language models (LLMs) can effectively recognize the student’s class actions and measure behavioral engagement based on identified actions, respectively. Accordingly, this paper addresses the following research questions:
@@
    Compared to traditional action classification models, can VLMs accurately classify students’ class actions?
@@
    Which is a more accurate depiction of students’ behavior: an action sequence or action histogram?
@@
  • We introduce a novel framework for measuring student behavioral engagement, which integrates VLM for action recognition and LLM for engagement classification. To overcome the lack of publicly available datasets for student actions and engagement, we perform few-shot tuning for VLM to improve student action recognition, and we leverage the LLM in zero-shot prompting for engagement classification.
  • Moreover, we construct a novel student behavior dataset with two components: 1) trimmed videos (13 categories, controlled environment) with annotated students’ actions and 2) untrimmed videos (3 lectures, 11 students) with annotated students’ engagement. The experts’ annotations of this dataset include engaged/disengaged labels for each 2-minute video and dense temporal action labels for each segment within them.
@@ 2 Related Work
2.1 Student Engagement Measurement
@@
The emerging and promising real-time engagement measurement approaches, enabled by advances in computer vision and machine learning, allows tracking a student’s engagement at fine-grained temporal resolutions on the order of seconds or minutes through discrete, objective indicators. Existing research of student behavior analysis in classrooms can be grouped into three categories: student action recognition, learning analytics, and engagement measurement. While student action recognition approaches focus on the student’s actions from trimmed videos or a single image, learning analytics approaches rely on facial expressions, student actions, and/or head gaze. For example, Sheng et al. [26] proposed an enhanced YOLOv8-based detector for localizing and classifying student behaviors from single classroom images. Their framework incorporates a Multi-Scale Large Kernel Convolution Module (MLKCM) to extract multiscale spatial features and a Progressive Feature Optimization Module (PFOM) that segments the channel dimension of the input feature map for improved feature refinement. Ahuja et al. [3] presented a classroom sensing system (EduSense) that provides a set of features, including detection of hand raise, hand on face, cross arms, smile, head orientation, and speech duration for both the teacher and the students. These features are extracted from students’ skeleton joints and facial landmarks due to the presence of much visual noise, occlusion in the classroom, and privacy preservation. Line et al.[22] recognized student behavior in the classroom based on upper body skeleton joints. They first extract students’ poses using bottom-up pose estimation (OpenPose) and remove the incorrect connection of students’ joints using YOLOv4 object detection, then generate a feature vector using normalized joint locations, joint distances, and bone angles to represent student posture and use a four-fully-connected-layer neural network to classify the student feature vector into four classes. The limitation of works [3, 22] is that student actions are recognized from single images rather than video segments, thereby ignoring temporal dynamics of the actions. Moreover, Line et al.[22] consider actions with different engagement levels, such as writing and playing with a phone, as one class (bowing).
@@
Student engagement studies, which employ student actions and/or head gaze to measure the student behavioral engagement, provide teachers with actionable insights. In [36], the authors proposed a framework that processes classroom videos by first estimating students’ 2D poses per frame and then applying a Temporal Action Detection (TAD) [17] to produce a temporal sequence of actions for each student. These per-student action sequences are subsequently fed into an LLM to generate behavioral reports. The TAD module was trained on the NTU RGB+D dataset [25], using action classes relevant to the classroom environment. However, their approach fails to provide explicit engagement labels and does not account for the peer-aware context when generating a textual behavioral report that includes each student’s actions analysis. Abdelkawy el al.[2] utilized a random forest classifier to detect whether the student is on-task or off-task based on a histogram of actions, together with head gaze. Specifically, each student’s 2-minute video was divided into overlapping segments using the sliding temporal window technique; the student skeleton sequence of each segment is assigned a predefined action label using 3D CNN model and then the histogram is created from these identified actions. Nevertheless, the student’s actions during the given period are not manually annotated. Consequently, the evaluation of the behavioral engagement classifier is still incomplete. Unlike these studies, we analyze the entire 2-minute video to get the student’s sequence of actions, which are used for engagement measurement. Moreover, we constructed a dataset that has dual expert annotations: engaged/disengaged for each 2-minute video and its corresponding frame-wise action labels to measure the performance of the engagement classifier. We also investigated the few-shot adaptation of the vision-language model for student action recognition, which learns to differentiate between action classes with a few training samples to overcome the challenge of the scarcity of annotated data required to model the diversity of student behaviors.
@@ 2.3 Large Language Models
Large language models (LLMs) are Transformer-based models with hundreds of billions of parameters, trained on large-scale text corpora (e.g., GPT-3 [9], LLaMA [13], Gemma [32], etc.). They exhibit generalization abilities, enabling them to perform tasks in zero-shot settings without requiring task-specific training examples. While several works have employed LLMs in education for multiple-choice question generation and assignment grading, Yu et al. [37] utilized an LLM to generate a pedagogical report based on the structured temporal action information for each student, including action categories and their timestamps. In contrast, our work uses an LLM to assign an engagement level to each student’s sequence of actions.
@@ 3.1 Overview of the proposed method
The proposed framework consists of three main steps: segment-wise action recognition, VLM action parsing, and engagement classification (Fig. 2). First, a vision-language model is adapted in a few-shot manner using a small set of labeled trimmed student action clips to improve recognition on continuous student videos. Second, the student’s untrimmed video is split into non-overlapping segments using a sliding-window technique, and the fine-tuned VLM assigns an action label to each segment. Consecutive identical predictions are merged to form an ordered action sequence that preserves both action identity and duration. Finally, this action sequence, along with the classroom context, is provided to an LLM to determine whether the student is engaged or disengaged.
Representing student’s behavior during a 2-minute interval as a temporal sequence rather than a histogram, where the duration of each action is accumulated, allows the model to capture not only how long each action occurs but also their ordering, which is shown to be crucial for interpreting engagement. For example, a student may still be deemed engaged if they take notes for the majority of the time period but play on their phone for one continuous 25-second block (e.g., at the beginning or end). However, a student is considered disengaged if they alternate between taking notes and making frequent quick phone checks because this pattern of frequent disruption shows an inability to maintain cognitive focus on the task. Therefore, the temporal sequence and context of actions more accurately reflect engagement than the total amount of time spent in each action (e.g., via a histogram). Furthermore, when representing student action as a sequence, together with peers’ action sequences, the LLM predictions are easier to interpret. Fig.3 shows examples of students’ actions aligned with classroom actions for engaged (a) and disengaged (b, c) students.
@@ 3.2 Actions dictionary
Because this study focuses on college students in introductory engineering STEM courses, the range of classroom actions from which behavioral engagement can be inferred is relatively limited. We adopted the classroom observation protocols in [20, 2], and identified the following 13 actions: Eating meal/snack, Writing on notebook/tablet, Typing on a laptop, Playing with mobile phone, Looking to the side/back, Looking down w/o reading/writing, Looking at laptop screen (not typing), Raising hand, Checking time, Reading, Drinking, Yawning , and Listening. Fig. 4 shows samples of such actions.